when Y=6

5y or y+19

Found 2 solutions by kryng88, JimMarshall:
Answer by kryng88(1)

y=6

5y
=5(6)
=30
or
y+19
=6+19
=25

Answer by JimMarshall(82)

y=6
5y
=5(6)
=30
or
y+19
=6+19
=25
Ans